### v1

[…]
Falis et al.
[…]
[$\ast$]Corresponding author: Matúš Falis, School of Informatics, The University of Edinburgh, 10 Crichton Street, Edinburgh EH8 9AB, United Kingdom; mfalis2@ed.ac.uk
Can GPT-3.5 Generate and Code Discharge Summaries?
[…]
(2023; 2023)
Can GPT-3.5 Generate and Code Discharge Summaries?
[…]
(2023; 2023)
keywords:
ICD Coding, Data Augmentation, Large Language Model, Clinical Text Generation, Evaluation by Clinicians
[…]
250 words in abstract; 3,929 words in main body; 2 figures (0 black and white, 2 colour); 4 tables; 34 references.
1 Abstract
Objective: To investigate GPT-3.5 in generating and coding medical documents with ICD-10 codes for data augmentation on low-resources labels.
Materials and Methods: Employing GPT-3.5 we generated and coded 9,606 discharge summaries based on lists of ICD-10 code descriptions of patients with infrequent (generation) codes within the MIMIC-IV dataset. Combined with the baseline training set, this formed an augmented training set. Neural coding models were trained on baseline and augmented data and evaluated on a MIMIC-IV test set. We report micro- and macro-F1 scores on the full codeset, generation codes, and their families. Weak Hierarchical Confusion Matrices were employed to determine within-family and outside-of-family coding errors in the latter codesets. The coding performance of GPT-3.5 was evaluated both on prompt-guided self-generated data and real MIMIC-IV data. Clinical professionals evaluated the clinical acceptability of the generated documents.
Results: Augmentation slightly hinders the overall performance of the models but improves performance for the generation candidate codes and their families, including one unseen in the baseline training data. Augmented models display lower out-of-family error rates. GPT-3.5 can identify ICD-10 codes by the prompted descriptions, but performs poorly on real data. Evaluators note the correctness of generated concepts while suffering in variety, supporting information, and narrative.
Discussion and Conclusion: GPT-3.5 alone is unsuitable for ICD-10 coding. Augmentation positively affects generation code families but mainly benefits codes with existing examples. Augmentation reduces out-of-family errors. Discharge summaries generated by GPT-3.5 state prompted concepts correctly but lack variety, and authenticity in narratives. They are unsuitable for clinical practice.
2 BACKGROUND AND SIGNIFICANCE
Large-Scale Multi-Labelled Text Classification (LMTC) tasks in Natural Language Processing (NLP) associate input documents with a set of output labels from a large label space, often hierarchically with a big-head long-tail distribution and data sparsity issues.
Medical document coding is the task of assigning structured codes from a medical ontology – e.g., the International Classification of Diseases (ICD) [^1] https://www.who.int/standards/classifications/classification-of-diseases – to clinical documents, a task performed by specially trained hospital staff. Coding consumes human resources that could be allocated to patient care. To ease this burden, research in machine learning and NLP cast medical document coding as an LMTC task [1].
In automatic ICD Coding, discharge summaries serve as input, yielding codes from a specified ICD version (e.g., ICD-10-CM) [^2] https://www.cdc.gov/nchs/icd/icd-10-cm.htm. ICD coding faces distribution challenges mirroring other LMTC tasks. Few common conditions (e.g., hypertension), contrast with many underrepresented or absent in corpora, such as MIMIC-IV [2]. Moreover, limited real-world data availability, often restricted for privacy reasons, compounds these challenges. However, modern deep learning ICD coding approaches (e.g., CAML [3], HLAN [4], RAC [5]) are data-driven, and adversely affected by data sparsity unless explicitly designed to handle label under-representation. Techniques such as auxiliary information [6, 7, 8, 9], or data augmentation and synthesis [10, 11, 12] attempt to mitigate these issues. ICD coding models with pre-trained encoders at best match the current state-of-the-art – usually involving domain-specific versions of BERT [13]. Large Language Models (LLMs) such as GPT-3 and its newer variants [14] (e.g., GPT-3.5) or Large Language Model Meta AI [LLaMA; 15] have recently displayed state-of-the-art performance on several tasks with emerging capabilities [16]. In the medical domain, notably Med-PaLM 2 [17] matching human performance in multiple-choice medical schools’ exams. While using models such as GPT-3.5 is problematic with real discharge summaries due to privacy issues, these models have the potential to aid in generating synthetic discharge summaries for training local models.
Recently, Large Language Models (LLMs), notably GPT-3.5, have become a new standard for advanced NLP tasks, especially ones reliant on understanding natural language. These models retain and apply background knowledge observed during training, yet can also produce fluent but inaccurate information (known as hallucination) [18]. Writing and coding discharge summaries requires extensive background knowledge making it of interest to explore an LLM’s capability of reading, coding, and generating discharge summaries.
An LLM capable of handling medical text could address the data sparsity issue by synthesising new data. This study aims to investigate the viability of GPT-3.5-generated medical documents for data augmentation in training local neural models and their credibility in clinical settings. We investigate GPT-3.5 given its performance in natural language understanding tasks requiring background knowledge, within an ethical experimental setting for clinical note data (by disabling content monitoring by the service provider). Furthermore, we explore GPT-3.5’s ability to code real discharge summaries and self-generated text.
Recent studies have prompted discussions on GPT’s utility in medicine, including applications in medical chatbots [19], or radiology [20]. Yeung et al. [21] compared the ChatGPT with a clinical GPT model [22] on generating patient vignettes. While research exists in generating data in low-resource settings in similar domains (e.g., law [23]), to the best of our knowledge, GPT’s performance in generating discharge summaries based on input conditions and its ability to perform ICD coding has not yet been reported.
OBJECTIVE
This study aims to assess GPT-3.5’s efficacy in the context of automated ICD-10 coding and investigate its viability as:
  • A data generator for ICD-10 coding, enhancing the training of local neural models by including GPT-3.5-generated discharge summaries, especially for rare labels;
  • An automated ICD-coding classifier, either using synthetic text with explicit code descriptions or real data as prompts;
  • A discharge summary generator focusing on producing clinically accurate and plausible synthetic data from expert perspectives.
MATERIALS AND METHODS
We queried GPT-3.5 (gpt-3.5-turbo) [^3] https://platform.openai.com/docs/models/gpt-3-5, from hereon referred to as G, through the OpenAI Python API [^4] https://platform.openai.com/docs/api-reference?lang=python to generate a patient discharge summary based on specific conditions and procedures represented by ICD-10-CM and ICD-10-PCS code descriptions from gold standard labels associated with MIMIC-IV discharge summaries (from the table hosp/d_icd_diagnoses.csv.gz). These produced label descriptions chosen to represent realistic combinations of conditions and procedures are not considered a part of the MIMIC-IV dataset. Note that sharing access to MIMIC data via any online API is prohibited [^5] https://physionet.org/news/post/415666Our method was consulted with and approved by PhysioNet, as we merely use the descriptions of attached codes.
While various dataset splits have been proposed since the release of coded discharge summaries in MIMIC-IV, we specifically chose not to follow the one proposed by Edin et al. [24] due to its exclusion of the long tail, which contrasts with our aim of addressing this aspect through generation techniques. Instead, we adopted the dataset split proposed by Nguyen et al. [25] that preserved the long-tail, aligning better with our focus. Nonetheless, we utilised the implementation of common ICD coding models produced by Edin et al. [24] for our analysis.
2.1 Label Selection
Candidate source documents for generation were selected from MIMIC-IV based on label populations in our selected split. We identified codes common across training, validation, and test sets, choosing those appearing up to 5 times in the training set, resulting in 195 unique codes (compared to 15,353 unique few-shot codes in training).
For these 195 codes we produced a list of codes belonging to their families (identified by the head of the code). Families with at least one relatively frequent code (population $>$ 100 in training) and at least one code exclusive to the test set (zero-shot) were retained leaving 16 families. We randomly selected 10 of these to generate from. For each family, we gathered its descendant codes appearing in the training set (139) and the zero-shot codes (16) totalling 155 codes. Of these, 114 codes had a population lower than 100 and are henceforth termed generation codes (list of codes available in Supplementary Material).
2.2 Preparation of Samples for Generation
In the Nguyen’s MIMIC-IV training set, we found documents with the 98 relevant few-shot codes (the 16 zero-shot were by-definition absent). Some documents contained multiple relevant codes. We have collected documents for each of the relevant codes and cloned them to bring their population up to 100. To reduce duplication, we have randomly dropped up to 5 of the assigned non-relevant labels within clones to create the new set of labels for generation (referred to as the silver standard).
We identified documents containing the siblings of the 16 zero-shot labels. A silver standard set was created for each of these documents substituting the sibling code with the zero-shot code (similar to the zero-shot approach in [10]). If multiple siblings were present, a random one was replaced with the zero-shot code.
This resulted in 9,606 input sets of labels – 6,779 unique and 2,827 duplicated.
2.3 Generation
Natural Language Generation is the task of producing natural language text based on a set of input data. We used the model “gpt-3.5-turbo-0613” given its wide recognition in the field of LLMs, relative cost-effectiveness, and time efficiency (compared to gpt-4). For distinct label sets, we utilised a temperature (parameter in the 0-1 range controlling randomness) of 0 to produce deterministic outputs. A temperature of 0.1 was employed for duplicates, allowing output variation.
Within the prompt (available in Supplementary Material), we state the task of writing a discharge summary for a patient with a list of standard descriptions of their conditions and procedures based on our silver standard. We also add further specifications:
  • Length of up to 4,000 words (following the maximum cutoff point in previous work [24]). The overall input and output token restriction of GPT-3.5 is 4,096 (inclusive of the prompt);
  • Inclusion of social and family history;
  • Anonymisation was required for personal and location data (due to uncertainty of the anonymity of the data used in training of GPT-3.5), maintaining numeric information when relevant despite potential removal in preprocessing;
  • Explicit ICD code mentions within the main text were to be avoided to prevent model association or potential errors;
  • Clear numeric values were preferred rather than ranges, especially for time-related codes;
  • Providing a specific concept for codes involving the umbrella term “other” encompassing a range of conditions;
  • Omission of the keyword “unspecified” present in standard descriptions opting for a more natural means of expression;
  • Coding of the discharge summary was to be positioned at the end in a regular pattern (codes in square brackets) for model coding assessment.
The generated documents were processed to find the discharge diagnoses and extract the assigned codes with a regular expression.
In total, we have generated 9,606 synthetic training documents. We removed all mentions of ICD-10 codes and pre-processed the documents the same as the baseline data. Then we merged these documents with the baseline training set (110,442 MIMIC-IV documents), forming the augmented training set (120,048 documents). The baseline and augmented settings used the same validation and test sets with 4,017 and 7,851 real documents, respectively.
Analysis between the generated text, and MIMIC-IV ICD-10-coded discharge summaries (the entire dataset, and the subset used for source label sets used in generation) showed differences in word count (per label and overall – Figures 1a and 1b) between synthetic and real data while retaining similar distributions of label counts (Figure 1c). The synthetic discharge summaries tend to be shorter which could be the result of GPT-3.5’s 4,096 token limit. Upon publication, we intend to make the synthetic discharge available.
Figure 1: A comparison between the discharge summary data in MIMIC-IV, the source data, and the generated discharge summaries. Subfigures 1a and 1b focus on the number of words in documents, indicating that the GPT-generated data generally contains fewer words overall and per assigned label compared to the real data from MIMIC-IV. Subfigure 1c demonstrates that, although there’s a variance in document size, the distribution of the number of labels per document remains relatively similar across the datasets.
2.4 Local Neural Models
Most recent LMTC neural architectures are encoder-decoder models whose encoder processes the input text to generate a latent representation.
Architectures using a non-BERT-like encoder (e.g., in CAML [3], LAAT [26], or Multi-Res CNN [27]) utilise non-contextual (e.g., Word2Vec [28]) word embeddings, while BERT([29])-like encoders (e.g., in PLM-ICD[30]) enable contextual token representation. The decoder determines a probability for each label based on the latent representation. A probability threshold determines positive predictions.
2.5 Evaluation
We have conducted four evaluation rounds:
  • Local Neural Model Evaluation: Assessing CAML, LAAT, and Multi-Res CNN models’ performance on Nguyen’s test set. Models were trained either solely on Nguyen’s training set or enhanced with data generated by G (the augmented training set).
  • GPT’s coding on real data: Evaluation of G’s coding ability on MIMIC-IV using Nguyen’s test set;
  • GPT’s coding on GPT-generated data: Evaluation of G’s coding ability on generated documents (with provided code descriptions in the prompt)
  • Acceptability of Generated Data in Clinical Practice: Reviewing G-generated discharge summaries by clinical professionals to gauge their suitability in clinical settings.
2.5.1 Local Neural Model Evaluation
We used the codebase and training procedure of Edin et al. [24] to train and evaluate CAML, LAAT, and Multi-Res CNN on the ICD coding task using Nguyen’s split. Each model is trained with 20 epochs on the training set. Throughout the training, we evaluated the model on the validation set, selecting the model with the highest mean average precision (MAP) from each run as the final model. Subsequently, this final model underwent evaluation on the test set. Additionally, we explored PLM-ICD, considered a state-of-the-art model for this task. However, its performance on the baseline and augmented data was notably lower than previously reported. This model has been reported to be unstable during training, which depends on random seeds. We decided against tuning the random seed, and only report the performance for CAML, LAAT, and Multi-Res CNN.
The models’ test performance was evaluated using standard information retrieval metrics commonly used in LMTC tasks – micro- and macro-averaged Precision, Recall, and F1 scores
Micro-averaging assigns equal weight to each prediction, favouring high-population classes (e.g., hypertension). Macro-averaging, in contrast, computes the performance for each unique label and averages across the label space, giving each label’s average result equal weight regardless of their population. This highlights poor performance in less common classes. Our primary evaluation metrics common with the majority of previous work are micro-F1 and macro-F1 scores. Metrics are further explained in Supplementary Material.
2.5.2 GPT’s coding on real clinical notes
We used the Azure AI Services API [^7] https://azure.microsoft.com/en-gb/products/ai-services to test GPT’s ability to assign diagnosis codes based on real clinical notes. The API returns a free text response which we further processed to retrieve the predicted ICD-10 codes. In postprocessing, we verify the response format. For correctly structured arrays of JSON objects we simply extract the predictions. For incorrectly structured outputs, we employ a regular expression pattern to extract all diagnoses and ICD code pairs. The result is a list of predicted diagnoses and corresponding ICD-10 codes for each clinical note. Figure 2 illustrates the API call workflow.
Figure 2: The workflow of the GPT-3.5 prediction. We used Azure AI Services API to query GPT-3.5 and we employed a postprocessing step to extract the predicted diagnoses and ICD-10 codes for each clinical note.
For reproducibility, we specified the model version and the API version as “gpt-3.5-turbo-0613” and “2023-03-15-preview”, respectively. All parameters were set to zero for deterministic responses from G, including temperature, top P, frequency penalty, and presence penalty. The system prompt directed G to act as a clinician assigning ICD-10 diagnosis codes to clinical notes, specifying the expected output format as JSON objects with keys “diagnosis” and “icd_code”. Table 1 presents all hyperparameter details. Our code implementation can be found in our Github repository  [^8] https://github.com/EdinburghClinicalNLP/chatgpt_icd_coding.
Parameter
Value
Model Name
gpt-3.5-turbo-0613
API Version
2023-03-15-preview
Temperature
0
Top P
0
Frequency Penalty
0
Presence Penalty
0
System Prompt
You are a clinician and your task is to assign diagnoses and the associated ICD-10 clinical codes to a given clinical note. You should output an array of JSON objects containing the diagnosis name and the ICD-10 code, structured as diagnosis and icd_code.
Table 1: Azure API call hyperparameters.
We opted out of human review of the data for two reasons. Firstly, the terms of the data use agreement of MIMIC-IV  [^9] https://physionet.org/news/post/415 did not grant us the authority to permit a third party to process the data for abuse detection. Secondly, we assessed the likelihood of harmful misuse to be low given the sensitive nature of the clinical notes.
In our evaluation of GPT’s performance, we have also employed hierarchical evaluation techniques – set-based hierarchical evaluation  [31] and Count-Preserving Hierarchical Evaluation (CoPHE) [32]. These metrics award partial credit to mispredicted labels by extending prediction and gold standard sets with their ancestor labels. While set-based evaluation ancestor labels track only presence of descendants, in CoPHE, ancestor labels link to the count of descendants, penalising over- and under-predictions within code families.
Comparing set-based and CoPHE results helps to evaluate the model’s tendency to over-/under-predict. A lower CoPHE score indicates this phenomenon. See Supplementary Materials for details on calculating hierarchical scores.
We utilise macro-averaged metrics from Weak Hierarchical Confusion Matrices (WHCM) [10] to summarise in-family versus out-of-family (OOF) prediction errors. These metrics are chosen to explore how expanding the population of codes within code families to a minimum of 100 instances impacts within-family performance. Within-family errors involve false positives that align with false negatives within the same family in the gold standard. On the other hand, an OOF error for a false negative in the gold standard lacks a false positive within the prediction set from the same family to match. Our primary goal in generating synthetic data is to reduce OOF errors, enhancing true positive predictions or ensuring mispredictions occur within-family.
2.5.3 GPT-3.5’s coding on Synthetic data
The prompt asked G to code the conditions and procedures mentioned in the document it generated. This experiment tested G’s ability to assign ICD-10 codes to concepts presented in their standard descriptions. Alongside this, the prompt required creating a patient’s social and family history, which might have led to the model introducing new conditions like substance abuse and potentially coding them, despite not being part of the initial prompt.
2.5.4 Acceptability of Generated Data in Clinical Practice
Four clinical professionals (co-authors SB, LD, MH, and RSP) assessed the quality of the generated data. As the data was generated based on labels associated with MIMIC-IV discharge summaries, this evaluation included both synthetic discharge summaries generated by G and discharge summaries from MIMIC-IV. The clinicians were presented with 20 discharge summaries – 10 synthetic and 10 real (based on whose adjusted gold standard the synthetic ones were generated).
Each discharge summary was assessed for:
  • Correctness – accuracy in describing patient conditions and procedures;
  • Informativeness – clarity and sense in supporting information (e.g., test results, medication suggestions);
  • Authenticity (patient) – whether such a patient could exist;
  • Authenticity (clinical scenario) – whether the hospital course was plausible as reported;
  • Acceptability – suitability of the document for clinical use.
Additionally, they separately evaluated correctness and informativeness for both non-low-resource and low-resource labels to gauge G’s ability to generate low-resource data. Scores from 1 (failure to perform) to 5 (perfect performance) were assigned to each metric, with accompanying comments justifying the score.
3 RESULTS
3.1 Local Neural Model Evaluation
We assessed performance across three code subsets: the entire codeset in MIMIC-IV ($overall$), a restricted set encompassing our candidate generation labels and their related codes ($f$), and a further subset ($f(gen)$) containing only the 114 low-population candidate generation labels. Results are shown in Table 2. Baseline results for CAML and LAAT align closely with prior findings by Nguyen et al. [25] for $overall$ metrics. While Nguyen et al. [25] did not report on Multi-Res CNN, its performance trends were similar to CAML and LAAT in comparison to Edin et al. [24] on non-filtered codesets. LAAT excels in micro-F1, Multi-Res CNN leads in macro-F1, and CAML generally lags behind. Baseline models outperform augmented ones in $overall$ micro-F1, a common observation when enhancing lower-resource label performance. Nonetheless, macro-F1 scores improved for two out of three models within the $overall$ codeset and for all models on $f$ and $f(gen)$. Multi-Res CNN and CAML macro-F1 scores display sizable relative improvement (26% and 78% respectively) in $f(gen)$.
Augmented models performed on par with or outperformed baseline models in micro-F1 scores for $f$ and $f(gen)$. Augmented Multi-Res CNN outperforms its baseline in micro-F1 for both $f$ and $f(gen)$, indicating benefits for the code family from augmenting less-populous members. Augmented LAAT shows improvement in macro-F1 in both $f$ and $f(gen)$ but lags in micro-F1
Comparing within-family and out-of-family errors (Table 3), augmented models generally exhibit fewer out-of-family errors on $f$ and $f(gen)$. An exception to this is augmented Multi-Res CNN in $f(gen)$, whose slight increase in OOF came with a sizable reduction in within-family error.
Unlike baseline models, augmented ones occasionally predicted codes absent from the original data, although incorrectly, except for one correctly predicted code (S02.63XA) by a Multi-Res CNN model trained on augmented data. While consistent enhancement in zero-shot code performance was not achieved through augmentation, the potential for improvement is evident.
Experiment $mi_{overall}$ $ma_{overall}$ $mi_{f}$ $ma_{f}$ $mi_{f(gen)}$ $ma_{f(gen)}$
CAML baseline 53.651 3.866 38.431 3.027 17.413 6.635
CAML augmented 53.544 3.899 38.410 3.780 20.680 11.858
LAAT baseline 57.285 6.181 43.586 4.959 26.790 14.478
LAAT augmented 57.176 6.091 43.364 5.381 25.696 14.978
Multi-Res CNN baseline 55.663 6.402 40.159 5.039 26.801 13.923
Multi-Res CNN augmented 54.694 6.464 42.688 5.846 30.390 17.681
Table 2: A comparison between local neural network models trained on baseline and augmented training sets is evaluated using micro- and macro-averaged $F1$ scores ($mi$ and $ma$ respectively) on three codesets – $overall$ on all codes present in MIMIC-IV; $f$ comprising all codes within the families we chose for generation; and $f(gen)$ corresponding to candidate codes used in generation with a population of at most 100 in the training set. The highest score in each metric for each model pair (baseline versus augmented) is highlighted in bold.
[…]
CAML baseline 66.525 25.045 83.809 9.831
CAML augmented 65.983 23.768 79.785 9.169
LAAT baseline 58.568 28.351 74.033 12.197
LAAT augmented 55.932 29.783 73.650 11.981
Multi-Res CNN baseline 52.720 32.406 69.676 15.242
Multi-Res CNN augmented 49.653 32.359 70.413 10.216
Table 3: Weak Hierarchical Confusion Matrix (WHCM) error rates of local neural network models trained on the original (baseline) training set versus on the training set augmented with synthetic data (augmented). Performance on the common test set is reported using the macro-averaged proportion of errors that were Out-of-Family (OOF) and within-family (In-Family). Errors are analysed for two codesets – $f$ on the codes within the families of codes used for generation; and $f(gen)$ corresponding to codes in $f$ with a population of at most 100 in the baseline training set. The best (lowest) error rate for each error type for each model pair (baseline versus augmented) is presented in bold.
3.2 GPT’s coding ability on real and synthetic data
We examined G’s ability to code real MIMIC-IV documents and generate coded documents with explicit code descriptions in the prompt. The results (Table 4) show that the performance on prompt-guided self-generated (synthetic) data resembles that of local models on the MIMIC-IV test set, not surpassing it. Hierarchical metrics show higher precision, recall, and consequently, F1-score in CoPHE compared to set-based hierarchical evaluation indicating errors coming from within-family misprediction, rather than incorrectly estimating the number of expected labels.
However, the performance on the MIMIC-IV test set is notably low, especially in precision. The improvement in the precision from leaf-only results to hierarchical is minimal. This implies that incorrect predictions were more likely to be out-of-family. Moreover, results on CoPHE are lower than on the set-based hierarchical evaluation indicating a tendency of the model to over-/under-predict within the scope of the family – an issue previously reported in local ICD coding models [32].
These results demonstrate that G can identify ICD-10 codes based on provided descriptions if presented within the prompts. Its performance when tasked with standard ICD coding without explicitly identified concepts or non-standard surface forms of the concepts significantly deteriorates.
Evaluation Set Leaf-Only Set-Based CoPHE
P R F1 P R F1 P R F1
Real 9.46 33.511 14.755 10.585 44.873 17.129 10.3 44.33 16.715
Synthetic 59.064 40.715 48.203 66.461 41.319 50.958 67.198 41.545 51.346
Table 4: Results of GPT’s coding ability on real and self-generated data.
3.3 Acceptability of Generated Data in Clinical Practice
We calculated the inter-evaluator agreement for the 7 metrics using Fleiss’ kappa ($\kappa$) [33]. As $\kappa$ is designed for categorical variables and does not fully capture ordinal scores, we present it alongside the mean scores for each metric ($\mu$). The evaluators’ agreement was poor ($\kappa<0$) in examples from MIMIC-IV for Correctness and Informativeness of non-low-resource codes, and the Acceptability of the discharge summaries. For the other metrics, a $\kappa>0$ was reached but never exceeded 0.4 (lower than moderate agreement). All mean scores are higher than 4. Hence, while the clinicians disagreed on the exact scores, they rated real discharge summaries positively. The disagreement may be due to clinicians being UK-based with significant differences in reporting style within the UK and the US (where MIMIC-IV is from).
For GPT-generated summaries, slight agreement was seen in acceptability, and fair agreement in the correctness of low-resource labels. All other metrics had poor agreement. Both correctness metrics scored above 4, with low-resource correctness surpassing 4.5—an encouraging outcome for our primary goal of generating low-resource code data. Mean informativeness in the low-resource scenario and authenticity of scores were at least 3. Once again, performance on the low-resource codes exceeded non-low-resource codes. Other metrics had $\mu$ scores above 2. Informativeness and authenticity for non-low-resource codes had poor agreement ($\kappa<0$), while acceptability had some agreement with the lowest mean score of 2.225. The full table of $\kappa$ and $\mu$ scores for the evaluation can be found in Supplementary Material.
While G exhibits a good level of correctness as affirmed by the clinicians, several challenges persist in the generation of natural-looking clinical notes:
3.3.1 GPT-3.5 tends to do verbatim reproductions of the prompted diagnoses list
G tends to generate clinical notes which list all concepts mentioned in the prompt. While an LLM following instructions is desirable behaviour, it is also important to exclude non-crucial details when tasked to generate a natural-looking clinical note. Real clinical notes often omit irrelevant normal findings and certain less-critical diagnoses within the context of an admission inferrable from supporting information such as medications and treatments for brevity. For instance, G unnecessarily noted a BMI within a normal range.
3.3.2 GPT-3.5 may phrase diagnoses in an unnatural manner
G may also express diagnoses in technical and verbatim manners, not typical in clinical text. For instance, G mentioned “anaemia, which was unspecified.” in the generated clinical note as it was prompted with “D64.9: Anemia, unspecified”. G also occasionally introduced vague phrases such as “geriatric team provided supportive care, including behavioural interventions and medication management” without further detail. This unnatural style easily differentiates synthetic clinical notes from real ones.
3.3.3 GPT-3.5 misses necessary supporting information
G tends to produce clinical notes lacking crucial supporting information. For instance, G mentioned “Following a traumatic event” without specifying the event, which is unacceptable in the clinical setting. G also mentioned the introduction of pleural intervention without additional detail or rationale. This omission limits the overall informativeness of the patient’s medical context, potentially hindering the notes’ usability for a comprehensive view.
3.3.4 GPT-3.5 may introduce spurious supporting information
The clinicians also observed that G sometimes introduces improbable but possible details. For instance, G overemphasised the significance of a patient’s anxiety disorder regarding an episode of syncope and a subsequent facial fracture, which the clinicians consider unlikely.
3.3.5 GPT-3.5 failed to present diagnoses as interconnected events
G also failed to link diagnoses as interconnected events, crucial for coherent patient summaries. The clinicians described G-generated clinical notes as collections of unrelated facts. For example, G listed multiple complications of Type 1 diabetes mellitus (i.e., “E10.43”, “E10.621”, “E10.628”, “E10.21”, “E10.319”) separately, failing to acknowledge their connectedness. Within the same note G failed to recognise the connection between vascular complications and diabetes, presenting them as independent facts. The lack of coherence between diagnoses may impede the plausibility of the clinical note and undermine the overall acceptability and usefulness of synthetic notes.
3.3.6 GPT-3.5 failed to prioritise and emphasise critical diagnoses
Furthermore, G failed at prioritising critical diagnoses, which undermined the authenticity of the portrayed scenario. For example, G presented multiple fractures alongside lesser issues, such as impacted ear wax, cataracts, and conjunctival haemorrhage. Hence, we concluded that G struggles to effectively convey the relative clinical significance of certain diagnoses.
4 DISCUSSION, CONCLUSION AND FUTURE STUDIES
In this work, we have investigated the capability of GPT-3.5’s potential in augmenting ICD-10 coding for local neural models in low-resource scenarios. While overall performance dipped with synthetic data augmentation, filtered codeset evaluation showed improvements, especially in advanced models like LAAT and Multi-Res CNN. Error analysis indicated augmented models made fewer out-of-family predictions, with some shift to within-family errors (closer to the correct answer). Augmentation showed promise in improving the prediction of generated codes and their siblings. Zero-shot labels did not consistently benefit from the augmentation, emphasising the need for real data in augmentation success. However, a zero-shot code learned from the synthetic data was predicted correctly.
In guided synthetic settings with ICD-10 descriptions, GPT-3.5 showed partial code identification ability displaying lesser over-/under-prediction tendencies than previously reported local models. It, however, struggled in the realistic scenario without in-prompt aid, performing below locally-trained models. Hence, it is unsuitable for deployment in a clinical setting.
Clinician-evaluated synthetic discharge summaries showed correctness in individual codes, yet lacked naturalness and coherence compared to real data, resulting in lower informativeness, authenticity, and acceptability scores. Synthetic summaries failed to represent holistic patient narratives or prioritise critical diagnoses.
One potential solution to generating synthetic discharge summaries involves restructuring the prompt to order diagnoses chronologically, providing their corresponding timestamps. This could guide LLMs in creating synthetic notes mirroring chronological progression of a patient’s medical journey, enhancing coherence and prioritisation.
Another promising solution is to retrieve real clinical notes as in-context learning examples to help guide the generation process [34] to aid LLMs in generating more realistic and coherent content. As this study focuses on evaluating LLMs’ existing capability, we opted to evaluate it in a zero-shot framework. Future work may explore this idea’s potential for generating more realistic-looking clinical notes.
5 Funding
This work is supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics. HD is supported by the Engineering and Physical Sciences Research Council (EPSRC, grant EP/V050869/1), Concur: Knowledge Base Construction and Curation. RSP is a fellow on the Multimorbidity Doctoral Training Programme for Health Professionals, which is supported by the Wellcome Trust [223499/Z/21/Z]. BA is supported by the Advanced Care Research Centre at the University of Edinburgh.
6 Competing Interests
We have identified no competing interests.
7 Author Contributions
MF proposed exploring data augmentation for ICD coding with GPT-3.5. and developed the main ideas of the publication with AG, HD, AB, and BA. MF, AG, HD, AB, and BA contributed to the overall experimental design. MF designed and performed the synthesis of new discharge summaries, and the evaluation of GPT’s ICD-10 coding performance on GPT-generated data. AG evaluated GPT’s ICD-10 coding performance on MIMIC-IV data. MF, AG, HD, and LD designed the evaluation of real versus generated discharge summaries by clinical staff. LD, SB, MH, RSP performed the clinical expert evaluation of the data. MF and AG analysed the results of the evaluation by clinicians. MF, AG, HD wrote the manuscript. All authors participated in editing the manuscript. BA and AB supervised the project and provided feedback and input on the experiments and analyses.